OpenStudy (safia21):

Find the area of a square with a perimeter of 36 inches.

OpenStudy (anonymous):

36 /4 = 9 9^2 = 81 81 inches squared

OpenStudy (lalaly):

36/4=9 9^2=81=area

OpenStudy (anonymous):

So if you know the perimeter is 36 inches and it's a square, it means each side must be 36/4 = 9 inches Therefore the area is 9 x 9 = 81 inches squared